OpenStudy (anonymous):

It takes 0.225 seconds/revolution for a body of mass of 25 grams to rotate a circle with a radius of 15 cm. What is the centripetal force acting on that body?

OpenStudy (welshfella):

force = m w r^2 F is in Newtons . mass in kgs and r in meters w = the angular velocity in radians / sec you can work this out from the given speed in revs / sec.

OpenStudy (anonymous):

how can I convert seconds/revolution to radians/sec?

OpenStudy (welshfella):

1 rev = 2 pi radians it travels 2 pi radians in 0.225 secs so in 1 second it travels 2 pi / 0.225 radians

OpenStudy (welshfella):

w = 8.89 pi rads/sec

OpenStudy (anonymous):

5x10^(-3) ?

OpenStudy (welshfella):

now you need to convert 25 gms to kgs and 15 cms to meters.

OpenStudy (welshfella):

what is 5x10^-3 ?

OpenStudy (anonymous):

the answer for the centripetal force?

OpenStudy (welshfella):

hmm..

OpenStudy (welshfella):

not quite - you need to multiply that by pi

OpenStudy (anonymous):

sorry i forgot to put pi, is it 0.0157?

OpenStudy (welshfella):

yes thats in newtons

OpenStudy (welshfella):

you can write it as 1.57 * 10^-2 N

OpenStudy (anonymous):

the answer is not in the choices :(

OpenStudy (anonymous):

i know other formulas maybe we can try that :)

OpenStudy (welshfella):

oh - maybe i've gone wrong somewhere - its been a while since i did these what are the choices?

OpenStudy (anonymous):

\[T = 2\pi(r) \over v\]

OpenStudy (anonymous):

0.731 N, 749.47 N, 187.368 dynes, 2,924 N

OpenStudy (jhannybean):

Isn't centripetal force \[\sf F_c = \frac{ mv^2}{r}\]

OpenStudy (welshfella):

yes where v = linear velocity

OpenStudy (anonymous):

yes @Jhannybean :)

OpenStudy (jhannybean):

since centripetal acceleration = \(\sf a_c = \dfrac{v^2}{r}\)

OpenStudy (jhannybean):

cool, I still remember \(\checkmark\) haha

OpenStudy (welshfella):

yea my memory has failed me here I should have checked

OpenStudy (welshfella):

i was too confident lol!

OpenStudy (anonymous):

\[T = \frac{ (2\pi(r)) }{ v }\]

OpenStudy (anonymous):